the third term of a G.P is 45 and the fifth term is 405. Find the G.P. if the common ratio r is positive
Answer by Boreal(15235)

a3=a1^r^2=45
a5=a3^r^2=a1r^4
a1=45/r^2
a1=405/r^4
45/r^2=405/r^4
405r^2=45r^4
9=r^2
r=3
Therefore, a2=15 and a1=5
5*3^n.
an=5*3^(n-1)
The series is 5,15,45,135,405...